OpenStudy (anonymous):

Find the derivative of x/((x^3)+2).

sam (.sam.):

Quotient rule, \[\frac{d}{d\, x}\left(\frac{f(x)}{g(x)}\right)\text{ = }\frac{g(x) f^{\prime }(x)-f(x) g^{\prime }(x)}{g(x)^2}\]

sam (.sam.):

\[\frac{d}{dx}=\frac{x^3-\left(3 x^2\right) x+2}{\left(x^3+2\right)^2}\]

OpenStudy (anonymous):

Thanks!